Constant Growth Model P=D1/(K-g)
Example 1 Example 2 Example 3 Example 4 Example 5 Example 6
D1 2.05 4.1 1.25 2.15 6.35 2.25
K 11% 15% 8% 18% 12% 13%
g 9% 10% 5% 10% 5% 8%
P= 0.041 0.205 0.0375 0.172 0.4445 0.1125
Calculating D1      = (1+g) * D0
D0 2.25 3.75 1.25 1.5 4.25 2
D1 2.43 4.125 1.3125 1.65 4.4625 2.16
K 13% 15% 8% 18% 12% 13%
g 8% 10% 5% 10% 5% 8%
P= 0.1215 0.20625 0.039375 0.132 0.312375 0.108
Zero Growth Model P=Do/K
D0 1.5 5 2
K 15% 5% 10%
P= 10 100 20
P/E Ratio Model Stock Price = Estimated EPS x Expected P/E
EPS growth 20% 15% 10% 25% 10%
Last Years EPS $           2.25 $           3.00 $           1.75 $           0.25 $           4.10
Expected P/E ratio 5 20 15 10 28
Estimated EPS ?

?

? ? ?
Stock Price= ? ? ? ? ?
CAPM RRR= RFR + (Beta x (Market Return - RFR))
Risk Free Rate 5.5% 3.5% 2.5% 4.5% 5.5%
BETA               1.30               2.00               1.25               2.50               0.75
Market Return 15% 13% 14% 11% 10%
Risk Free Rate 5.5% 3.5% 2.5% 4.5% 5.5%
RRR 17.85% 22.50% 16.88% 20.75% 17.13%

Solve for estimated EPS and stock price.
Answer #1

example 1

estimated EPS=2.25×1.20=2.7

estimated stock price

=(2.25* 1.20)×5=13.5

example 2

Estimated eps=3×1.15=3.45

estimated stock price=3.45×20=69

example 3

estimated EPS=1.75×1.1=1.925

Estimated stock price=1.925×15=28.875

example 4

estimated EPS=0.25×1.25=0.3125

estimated stock price=0.3125×10=3.125

example 5

estimated stock price=4.10×1.1=4.51

estimated stock price=4.51×28=126.28.